# Mesoscopic Modeling of High-Density Carbon Nanotube Films for Memristive Device Applications

### v1

@@ 2.5 Electrical current calculation
To compute the electrical current flowing through the carbon nanotube networks, we employ a nodal analysis approach, adapted from the work of Jin et al. 49. Each CNT bead is treated as a node, and electrical connections between nodes are encoded into a sparse conductance matrix . The overall structure is modeled as a resistor network where both intratube and intertube conductances are explicitly accounted for. We define a local intratube conductance between adjacent beads of the same nanotube, and an intertube conductance between connected segments belonging to different CNTs.
Defect-free CNTs behave as one-dimensional ballistic conductors with a mean free path exceeding for the bias regime considered here 50 and a conductivity of independently of the tube length 51, 52, where is the conductance quantum. The intratube conductance is therefore given by:
@@
Based on the model of Durrant et al. combining density functional tight-binding (DFTB) theory and the non-equilibrium Green functions (NEGF) method 52, we estimate the average conductance at the junctions between CNTs for both chiralities as:
We note that tubes with larger diameters have lower junction conductance, an effect attributed to the reduced overlap between the delocalized conduction channels and the local contact area: as the tube circumference increases, the fraction of the electronic wave function contributing to the actual contact region decreases 52, 53.
@@
We considered two cases: (i) all CNTs are metallic, and (ii) since many CNT chiralities are semiconducting, only one third of the tubes are randomly selected as being metallic, while the remaining semiconducting ones are assumed to be non-conductive under the 0.5 V bias applied here. In the present model, the electrical conductance of each connection is treated as Ohmic and independent of the carrier statistics, such that both the fully metallic and partially metallic networks follow the same linear transport scheme. In reality, when only one third of the CNTs are metallic, the system would likely operate in a mixed regime where quantum conduction along metallic paths coexists with thermally activated hopping across semiconducting barriers 54, 55. Such effects are not explicitly captured here, but are expected to primarily rescale the overall conductance without altering the qualitative trends observed between structural descriptors and transport efficiency.
After solving the linear system given by Eq. (11), we extract the set of bonds that carry a significant current and reconstruct the connected subgraph linking the source and drain electrodes in order to identify the available dominant conduction pathways within the network. The possible contributions of aC particles to the total film conductivity are neglected in the present analysis.
3 Results and Discussion
@@
Due to the inherent variability in the initial random configurations and the necessity of applying a slightly different compression protocol for structures #30 and #31 (combining high-density, intermediate amorphous carbon content, and large-diameter tubes), the quantitative and physical interpretation of the descriptor correlations remains exploratory. A statistically more robust analysis would require multiple realizations per condition and a more uniform compression protocol across all structures. Nevertheless, the consistent qualitative trends observed across the dataset, in particular for buckling and bundling, indicate that the selected set of descriptors captures relevant topological and geometrical features of the CNT networks and provides a useful basis for future more systematic studies. We note that we were unable to identify the specific reason why the structures #30 and #31 developed a noticeably different morphology compared to the other aC-containing films and consider them representative structures.
@@
Low-density films composed of short CNTs (15 or 40 nm in length; #1#4 and #17#20) tend to exhibit large voids and pronounced bundle formation, reflected in high values of the bundling factor and average bundle size , while showing no detectable buckling (). According to the stability criterion proposed by Volkov et al. 34, which identify stable networks as those that avoid fragmentation into isolated and weakly connected bundles, these systems would be considered unstable. These configurations consistently show poor effective connectivity () and low electrical conductivity, particularly for high value of . For low-density films with 100-nm long tubes we observe a reduction in bundling and a concurrent increase in bending, curvature and current for the (16,0) chirality. These trends remain nearly unchanged for the large-diameter (32,0) tubes, showing that thinner CNTs are more prone to mechanical deformation under compression, which in turn improves intertube contacts and electrical conduction.
@@
In contrast, high-density structures exhibit significantly fewer voids, especially in the four-layer configurations and for the (32,0) chirality. This trend is not surprising. On the one hand, (32,0) films at 0.6 gcm-3 are already close to their close-packed density, whereas the corresponding value for (16,0) films is approximately 1.5 times higher, leaving more space for void formation. On the other hand, since the position of the voids within each layer is random, the voids do not align when the layers are stacked together. We note that without lateral confinement along the -direction several of these high-density films would expand and adopt a lower-density structure. If stability were judged on the basis of the ability of the films to maintain their morphology in vacuum, some high-density films would likely be considered unstable. However, in the capacitor geometry considered here, the lateral confinement is intrinsic, and the film density remains constant throughout. For these dense morphologies, layering systematically reduces electrical conductivity, whereas this trend is not observed for their low-density counterparts. The role of amorphous carbon appears to be more versatile, with its influence sensitively depending on the combined effects of chirality, mass fraction, and compression protocol. These observations highlight the complex interplay between structural parameters such as chirality, tube length, density, and amorphous carbon content that collectively dictate the final morphology and transport performance of the network.
@@
To quantitatively assess the relationships between structural descriptors and transport properties, we first computed the Pearson correlation matrix 56 across all available parameters, as shown in Fig. 4(a). Several strong correlations emerge, providing insight into how geometry and topology affect the electrical performance of CNT networks. Local geometric distortions, captured by the curvature factor and buckling factor , are among the most influential descriptors. These two quantities are strongly correlated (), as both reflect geometric distortions at the nanotube segment level. In particular, both and show a strong positive correlation with the total current ( and , respectively) and even more with the current when only one third of the tubes are considered metallic ( and ). These results suggest that local mechanical distortions facilitate charge transport, probably by enhancing intertube overlap and contact and/or by reducing the number of junctions along the main conductive paths.
@@
In contrast, the bundling factor shows a strong negative correlation with () and with (), indicating that highly bundled networks tend to conduct less efficiently. Similar trends have been widely reported in low-density, transparent CNT films, where increasing bundle size reduces the number of effective intertube junctions and limits charge percolation 57, 58, 59, 60, 55, 61. The geometric arguments made in these papers are consistent with the behavior we observe in our high-density films.
@@
The effective connectivity exhibits a moderate positive correlation with both currents ( with and with ), indicating that global network connectivity contributes to transport, though to a lesser extent than local geometric factors such as curvature and buckling, which display stronger correlations. This highlights the combined influence of local and global structural features on charge conduction. Interestingly, the minimum number of connections to the electrodes shows only a weak correlation ( and , respectively), which does not rule out a more significant role of the electrodes under cycling or nonequilibrium conditions.
@@
Global structural parameters also influence connectivity. Tube length is positively correlated with curvature (), buckling (), (), and both currents ( for and for ). Longer tubes are more likely to establish multiple contacts and undergo elastic deformation, thereby improving network connectivity. Film density correlates moderately with , , and (), as well as with total current (), consistent with the expected increase in intertube interactions in denser networks. In contrast, the number of layers is anticorrelated with connectivity and transport ( with currents, with ), likely reflecting the tendency of multilayer films to introduce additional junctions along percolation paths. Finally, the tube orientation angle and its dispersion are highly correlated with each other (), but show negligible correlation with transport properties or connectivity descriptors, indicating that the macroscopic angular disorder does not significantly affect global transport in the present dataset. However, the effects of local alignment or anisotropic percolation pathways may still play a role beyond the scope of the linear correlation analysis presented here.
@@
The role of amorphous carbon is complex and highly structure-dependent. Although it correlates positively with curvature () and buckling (), its correlation with transport is less clear. No monotonic trend is observed between the aC content and the current. This likely reflects the inhomogeneous spatial distribution of aC and the non-uniform generation protocols. In particular, structures with chirality (32,0) and 20-30% aC content (#30 and #31) required additional pressure-compression cycles, leading to the formation of large aC clusters and modified bundle morphologies. Such artifacts likely affect transport independently of the nominal aC percentage. For (16,0) structures (#13-#16), aC consistently aligns along CNTs, whereas (32,0) structures exhibit more diverse behaviors: junction clustering at 10% aC (#29), plaque formation at 20-30% (#30-#31), and tube alignment at 40% (#32). These variabilities highlight the sensitivity of aC morphology to both diameter and processing history. Consequently, conclusions regarding the role of aC should be drawn with caution. A more rigorous assessment would require the controlled structure generation and multiple realizations per configuration to statistically capture the influence of aC on film morphology and transport. These observations also suggest that it should be possible to tune the morphology of aC-containing films through careful control of the construction protocol, enabling the modeling of different experimental scenarios depending on the targeted application or physical conditions. Finally, we note that the aC content is generally expected to reduce the conductivity of the film by introducing additional carrier scattering 62, an effect distinct from the morphological influence discussed here.
@@
To test the robustness of the trends identified with the Pearson correlation matrix, we also calculated the Spearman rank correlation matrix 63, as shown in Fig. 4(b), which captures monotonic and possibly nonlinear relationships between variables. Overall, the Spearman analysis confirms most of the trends previously observed, particularly the strong associations between curvature, buckling, and transport properties. Some correlations involving the amorphous carbon content were slightly reduced, suggesting that its influence on transport is not strictly linear. In contrast, the correlations between angular descriptors ( and ) and current values were more pronounced by the Spearman analysis, hinting at potential nonlinear effects of tube orientation that are not well captured by Pearson correlation. This observation suggests that local alignment may influence transport pathways in a more subtle, nonlinear manner. The overall consistency between the two analyses reinforces the relevance of the identified descriptors, while also emphasizing the importance of considering nonlinear interactions in future modeling efforts.
@@
Finally, we performed a principal component analysis (PCA) 64, 65, 66, 67, 68 on the entire set of normalized structural descriptors, including geometrical and topological quantities. Electrical currents and were not included as input variables. However, their subsequent projection into the PCA space reveals how they correlate with the underlying structural patterns. The results for the most significant principal components are shown in Table S2. This dimensionality reduction reveals dominant patterns of variation across the dataset and facilitates the interpretation of the underlying structure-property relationships. The first principal component (PC1) alone captures 38.9% of the total variance, while PC2 accounts for an additional 19.3%. Together, the first three components explain more than 71% of the variance.
@@
In the PCA projection along the PC1-PC2 plane, shown in Fig. 5, individual structures can be grouped into clusters, in particular according to tubes length, with tightly grouped configurations observed for aC-containing structures (cluster A). Apart from structures #30 and #31, networks containing aC are the most conductive. Transport vectors (, ), when added as supplementary arrows in the correlation plot, point strongly along the negative PC1 direction, confirming that well-connected, low-bundling structures with high curvature and buckling consistently produce higher current output. In contrast, structures with weak connectivity and strong bundling, characterized by positive PC1 scores, fall into the low-performance regime. These findings highlight the value of PCA not only as a tool for reducing dimensionality but also as a means to identify collective structural signatures that govern macroscopic transport.
4 Conclusion
@@
We have presented a mesoscopic modeling framework for simulating high-density carbon nanotube networks with different structural morphologies and amorphous carbon content and to analyze their structural characteristics and conductivity in device applications. A diverse set of 32 CNT films was constructed, systematically varying chirality, tube length, network density, and aC content. The resulting structures were characterized by a set of geometrical and topological descriptors, and electrical current through the films was calculated with a nodal analysis framework. Additionally, we investigated the role of layering by comparing single- and multi-layer films, and found that its impact on connectivity and transport depends significantly on structural context, including chirality and tube length. To account for amorphous carbon, we introduced a simple yet effective modeling strategy enabling its integration into CNT films. Our analysis reveals that aC content can significantly alter the local morphology and affect transport, although in a nonmonotonic and configuration-dependent manner.

### v2

@@ 2.5 Electrical current calculation
To compute the electrical current flowing through the carbon nanotube networks, we employ a nodal analysis approach, adapted from the work of Jin et al. 23. Each CNT bead is treated as a node, and electrical connections between nodes are encoded into a sparse conductance matrix . The overall structure is modeled as a resistor network where both intra-tube and inter-tube conductances are explicitly accounted for. We define a local intra-tube conductance between adjacent beads of the same nanotube, and an inter-tube conductance between connected segments belonging to different CNTs.
Defect-free CNTs behave as one-dimensional ballistic conductors with a mean free path that exceeds for the bias regime considered here 38 and a conductivity of independently of the tube length 14, 11, where is the quantum conductance. The intra-tube conductance is therefore given by:
@@
Based on the model of Durrant et al. combining density functional tight-binding (DFTB) theory and the non-equilibrium Green functions (NEGF) method 11, we estimate the average conductance at the junctions between CNTs for both chiralities as:
We note that tubes with larger diameters have lower junction conductance, an effect attributed to the reduced overlap between the delocalized conduction channels and the local contact area: as the tube circumference increases, the fraction of the electronic wave function contributing to the actual contact region decreases 11, 7.
@@
We considered two cases: (i) all CNTs are metallic, and (ii) since many CNT chiralities are semiconducting, only one third of the tubes are randomly selected as being metallic, while the remaining semiconducting ones are assumed to be non-conductive under the 0.5 V bias applied here. While specific chiralities are used to define CNT geometries in the film construction, the subsequent electrical transport calculations are independent of chirality and depend only on the nanotube diameter. In the present model, the electrical conductance of each connection is treated as Ohmic and independent of the carrier statistics, so that the fully metallic and partially metallic networks follow the same linear transport scheme. In reality, when only one third of the CNTs are metallic, the system would likely operate in a mixed regime where quantum conduction along metallic paths coexists with thermally activated hopping across semiconducting barriers 65, 22. Such effects are not explicitly captured here, but are expected to primarily rescale the overall conductance without altering the qualitative trends observed between structural descriptors and transport efficiency.
After solving the linear system given by Eq. (11), we extract the set of bonds that carry a significant current and reconstruct the connected subgraph linking the source and drain electrodes to identify the available dominant conduction pathways within the network. We treat aC particles as electrically inactive inclusions, affecting the network morphology but not contributing directly to the total film conductivity, under the assumption that their conductivity is much lower than that of CNTs; their inclusion as conductive elements can nevertheless be straightforwardly implemented within the same framework in future work.
3 Results of Calculations
3.1 Film structure
@@
Due to the inherent variability in the initial random configurations and the necessity of applying a slightly different compression protocol for structures #30 and #31 (combining high-density, intermediate amorphous carbon content and large-diameter tubes), the quantitative and physical interpretation of the descriptor correlations remains exploratory. A statistically more robust analysis would require multiple realizations per condition and a more uniform compression protocol across all structures. Nevertheless, the consistent qualitative trends observed across the dataset, in particular for buckling and bundling, indicate that the selected set of descriptors captures relevant topological and geometrical features of the CNT networks and provides a useful basis for future more systematic studies. Although we were unable to identify the specific reason why the structures #30 and #31 developed a different morphology compared to the other aC-containing films, we consider them representative structures.
@@
Low-density films composed of short CNTs (15 or 40 nm in length; #1#4 and #17#20) tend to exhibit large voids and pronounced bundle formation, reflected in high values of the bundling factor and average bundle size , while showing no detectable buckling (). According to the stability criterion proposed by Volkov et al. 52, which identifies stable networks as those that avoid fragmentation into isolated and weakly connected bundles, these systems would be considered unstable. These configurations consistently show poor effective connectivity () and low electrical conductivity, particularly for high value of . For low-density films with 100-nm long tubes we observe a reduction in bundling and a concurrent increase in bending, curvature, and current for the (16,0) chirality. These trends remain nearly unchanged for the large-diameter (32,0) tubes, showing that thinner CNTs are more prone to mechanical deformation under compression, which in turn improves inter-tube contacts and electrical conduction.
In contrast, high-density structures exhibit significantly fewer voids, especially in the four-layer configurations and for the (32,0) chirality. This trend is not surprising. On the one hand, (32,0) films at 0.6 gcm-3 are already close to their close-packed density, whereas the corresponding value for (16,0) films is approximately 1.5 times higher, leaving more space for void formation. However, since the position of the voids within each layer is random, the voids do not align when the layers are stacked together.
@@
We note that without confinement in the -direction, several of these high-density films would expand and adopt a lower-density structure. If film stability were judged on the basis of their ability to maintain their morphology in vacuum, some high-density films would likely be considered unstable. However, in the capacitor geometry considered here, the lateral confinement is intrinsic and the film density remains constant throughout. For these dense morphologies, layering systematically reduces electrical conductivity, whereas this trend is not observed for their low-density counterparts. The role of amorphous carbon appears to be more versatile, with its influence sensitively depending on the combined effects of chirality, mass fraction, and compression protocol. These observations highlight the complex interplay between structural parameters, such as chirality, tube length, density, and amorphous carbon content, which collectively dictate the final morphology and transport performance of the network.
3.2 Correlation between structure and electrical transport
@@
To quantitatively assess the relationships between structural descriptors and transport properties, we first computed the Pearson correlation matrix 5 across all available parameters, as shown in Fig. 4(a). Several strong correlations emerge, providing insight into how geometry and topology affect the electrical performance of CNT networks. Local geometric distortions, captured by the curvature factor and buckling factor , are among the most influential descriptors. These two quantities are strongly correlated (), as both reflect geometric distortions at the nanotube segment level. In particular, both and show a strong positive correlation with the total current ( and , respectively) and even more with the current when only one third of the tubes are considered metallic ( and ). These results suggest that local mechanical distortions facilitate charge transport, probably by enhancing inter-tube overlap and contact and/or by reducing the number of junctions along the main conductive paths.
@@
In contrast, the bundling factor shows a strong negative correlation with () and with (), indicating that highly bundled networks tend to conduct less efficiently. Similar trends have been widely reported in low-density transparent CNT films, where increasing the size of the bundle reduces the number of effective inter-tube junctions and limits charge percolation 20, 33, 45, 35, 22, 44. The geometric arguments made in these papers are consistent with the behavior we observe in our high-density films.
@@
Effective connectivity exhibits a moderate positive correlation with both currents ( with and with ), indicating that global network connectivity contributes to transport, though to a lesser extent than local geometric factors such as curvature and buckling, which show stronger correlations. This highlights the combined influence of local and global structural features on charge conduction. Interestingly, the minimum number of connections to the electrodes shows only a weak correlation ( and , respectively), which does not rule out a more significant role of the electrodes under cycling or nonequilibrium conditions.
@@
Global structural parameters also influence connectivity. The length of the tube is positively correlated with curvature (), buckling (), (), and both currents ( for and for ). Longer tubes are more likely to establish multiple contacts and undergo elastic deformation, thereby improving network connectivity. Film density correlates moderately with , , and (), as well as with total current (), consistent with the expected increase in inter-tube interactions in denser networks. In contrast, the number of layers is anticorrelated with connectivity and transport ( with currents, with ), likely reflecting the tendency of multilayer films to introduce additional junctions along percolation paths. Finally, the tube orientation angle and its dispersion are highly correlated with each other (), but show negligible correlation with transport properties or connectivity descriptors, indicating that the macroscopic angular disorder does not significantly affect global transport in the present dataset. However, the effects of local alignment or anisotropic percolation pathways may still play a role beyond the scope of the linear correlation analysis presented here.
@@
The role of amorphous carbon is more complex and highly structure-dependent. Although it correlates positively with curvature () and buckling (), its correlation with transport is less clear. No monotonic trend is observed between the aC content and the current. This likely reflects the inhomogeneous spatial distribution of aC and the non-uniform generation protocols. In particular, structures with chirality (32,0) and 20-30% aC content (#30 and #31) required additional pressure-compression cycles, leading to the formation of large aC clusters and modified bundle morphologies. Such artifacts likely affect transport independently of the nominal aC percentage. For (16,0) structures (#13-#16), aC consistently aligns along CNTs, while (32,0) structures exhibit more diverse behaviors: junction clustering at 10% aC (#29), plaque formation at 20-30% (#30-#31), and tube alignment at 40% (#32). These variabilities highlight the sensitivity of aC morphology to both diameter and processing history. Consequently, conclusions about the role of aC should be drawn with caution. A more rigorous assessment would require the controlled structure generation and multiple realizations per configuration to statistically capture the influence of aC on film morphology and transport. These observations also suggest that it should be possible to tune the morphology of aC-containing films through careful control of the construction protocol, enabling modeling of different experimental scenarios depending on the targeted application or physical conditions. Finally, we note that the aC content is generally expected to reduce the conductivity of the film by introducing additional carrier scattering 15, an effect distinct from the morphological influence discussed here.
3.3 Trend analysis
@@
To test the robustness of the trends identified with the Pearson correlation matrix, we also calculated the Spearman rank correlation matrix 19, as shown in Fig. 4(b), which captures monotonic and possibly nonlinear relationships between variables. Overall, the Spearman analysis confirms most of the trends previously observed, particularly the strong associations between curvature, buckling, and transport properties. Some correlations involving the amorphous carbon content were slightly reduced, suggesting that its influence on transport is not strictly linear. In contrast, the correlations between angular descriptors ( and ) and current values were more pronounced by the Spearman analysis, hinting at potential nonlinear effects of tube orientation that are not well captured by Pearson correlation. This observation suggests that local alignment may influence transport pathways in a more subtle, nonlinear manner. The overall consistency between the two analyses reinforces the relevance of the identified descriptors, while also emphasizing the importance of considering nonlinear interactions in future modeling efforts. While pairwise Pearson and Spearman correlation matrices provide valuable insight into individual descriptor–transport relationships, they do not capture the collective influence of multiple correlated descriptors. To address this, we next employ principal component analysis (PCA) 62, 1, 24, 48, 17.
@@
Given the limited number of simulated structures, we performed a PCA analysis as an exploratory tool on the entire set of normalized structural descriptors, including geometrical and topological quantities. Electrical currents and were not included as input variables. However, their subsequent projection into the PCA space reveals how they correlate with the underlying structural patterns. The results for the most significant principal components are shown in Table S2. This dimensionality reduction reveals dominant patterns of variation across the dataset and facilitates the interpretation of the underlying structure-property relationships. The first principal component (PC1) alone captures 38.9% of the total variance, while PC2 accounts for an additional 19.3%. Together, the first three components explain more than 71% of the variance.
@@
In the PCA projection along the PC1-PC2 plane, shown in Fig. 5, individual structures can be grouped into clusters, in particular according to the length of tubes, with tightly grouped configurations observed for aC-containing structures (cluster A). Apart from structures #30 and #31, networks containing aC are the most conductive. Transport vectors (, ), when added as supplementary arrows in the correlation plot, point strongly along the negative PC1 direction, confirming that well-connected, low-bundling structures with high curvature and buckling consistently produce higher current output. In contrast, structures with weak connectivity and strong bundling, characterized by positive PC1 scores, fall into the low-performance regime. These findings highlight the value of PCA not only as a tool for reducing dimensionality but also as a means to identify collective structural signatures that govern macroscopic transport.
4 Discussion
@@
We have presented a mesoscopic modeling framework for simulating high-density CNT networks with different structural morphologies and amorphous carbon content and analyzed their structural characteristics and electrical transport properties. A diverse set of 32 CNT films was constructed, systematically varying chirality, tube length, network density, and aC content. The resulting structures were characterized by a set of geometrical and topological descriptors, and electrical current through the films was calculated with a nodal analysis framework. Additionally, we investigated the role of layering by comparing single- and multi-layer films and found that its impact on connectivity and transport depends significantly on the structural context, including chirality and tube length. To account for amorphous carbon, we introduced a simple yet effective modeling strategy enabling its integration into CNT films. Our analysis reveals that the aC content can significantly alter the local morphology and thereby influence electrical transport, although in a nonmonotonic and configuration-dependent manner. However, the presented modeling framework relies on a set of simplifications that define the range of applicability of the results reported here.
First, the simulated networks are composed exclusively of single-walled carbon nanotubes (SWCNTs), as the current coarse-grained model does not include an explicit representation of multiple concentric CNT walls or their associated mechanical response. As a result, the quantitative magnitude of curvature- and buckling-related effects reported in this work reflects the mechanical compliance of SWCNT-based networks. While the underlying mesoscale framework remains applicable in principle, extending the present analysis to multi-walled CNT (MWCNT) networks would require the explicit inclusion of MWCNT-specific mechanical properties, which may modify the relative weight of deformation-related descriptors without altering their qualitative role.
Secondly, amorphous carbon (aC) particles are treated here as electrically inactive inclusions that influence the network morphology but do not contribute directly to electrical transport. This modeling choice allows us to isolate the impact of aC on CNT packing, bundling, and junction statistics. In the generated structures, aC particles are predominantly attached to CNT sidewalls rather than occupying inter-tube junction regions, so that the dominant current pathways remain CNT-mediated. Neglecting the conductive contribution of aC may therefore affect the quantitative distribution of current and the detailed topology of percolating paths in systems where aC participates actively in transport. However, given their size and spatial arrangement, aC-mediated pathways would typically involve a larger number of junctions, making them less efficient than CNT-dominated routes within the present structural regime. This modeling choice is therefore physically justified for the configurations investigated here.
Thirdly, the structural descriptors introduced in this study are analyzed within a static framework and are used to characterize baseline structure-transport relationships. For the applications that involve dynamic, field-driven structural evolution, the descriptors identified here provide a natural basis for future extensions. For instance, field-induced filament formation or junction reconfiguration in RRAM would be expected to manifest as changes in effective connectivity , in the minimum number of junctions along dominant current paths, or in the bundling factor . In this context, the present results identify which descriptors are most relevant for transport and therefore which quantities are likely to be most sensitive to electrically driven structural evolution, without explicitly modeling switching dynamics.
Finally, the quantitative trends reported in this work pertain to a specific range of CNT lengths, network densities, and aC contents motivated by experimentally fabricated dense CNT films. In networks composed of substantially longer CNTs, increased aspect ratios would enhance bridging capabilities and increase the number of junctions per tube, potentially modifying the relative importance of local descriptors such as bundling. Extending the present analysis to such regimes would require additional global topological descriptors such as entanglement and tortuosity which are expected to become increasingly relevant. In particular, while increased entanglement or tortuosity may counterbalance the negative impact of bundling by improving global connectivity, the local effect of bundling in increasing the number of effective inter-tube junctions along a given pathway is expected to remain qualitatively unchanged. We believe that the underlying mesoscale modeling approach and the associated descriptor-based analysis presented here are general and can be readily applied to a broader class of CNT-based networks and composite architectures.
5 Conclusions